OpenStudy (anonymous):

Solve for x: 25x^2 − 36 = 0 x = 6, 5 x = −6, −5 x = 36 over 25, x = − 36 over 25 x = 6 over 5, x = − 6 over 5

OpenStudy (solomonzelman):

\(\large\color{black}{ \rm 25x^2 − 36 = 0 }\) \(\large\color{black}{ \rm 25x^2 = 36 }\) \(\large\color{black}{ \rm 5^2x^2 = 36 }\) \(\large\color{black}{ \rm (5x)^2 = 36 }\) \(\large\color{black}{ \rm (5x)^2 = (±6)^2 }\) \(\large\color{black}{ \rm 5x = ±6 }\) \(\large\color{black}{ \rm x = ±\frac{6}{5} }\)

OpenStudy (anonymous):

thank you

OpenStudy (solomonzelman):

Anytime !